OpenStudy (anonymous):

RATIONAL EXPRESSION HELP X/X^2-1+6X/X-X^2

OpenStudy (dumbcow):

Factor denominators \[\frac{x}{(x+1)(x-1)} + \frac{6x}{x(1-x)}\] cancel x"s .... pull out a neg from 2nd denominator \[\frac{x}{(x+1)(x-1)} - \frac{6}{x-1}\] Notice both denominators share the factor (x-1) To make denominators the same multiply 2nd fraction by (x+1)/(x+1) \[\frac{x}{(x+1)(x-1)} - \frac{6(x+1)}{(x+1)(x-1)}\] Now combine fractions... add like terms on top \[\frac{x -6x-6}{(x+1)(x-1)}\] \[\frac{-5x-6}{(x+1)(x-1)}\]